OpenStudy (anonymous):

Here is an addition sentence. (3 + 1) + 11 = 15 Choose an equivalent addition sentence that shows the associative property of addition.

OpenStudy (neer2890):

(2+2)+11=2+(2+11) 15=15

OpenStudy (anonymous):

them is not one of the answer choices.

OpenStudy (anonymous):

these are the answer choices (3 + 1) + 11 = 15 3 + (1 + 11) = 15 15 = (3 + 1) + 11 3 + 1 + 11 = 15

OpenStudy (neer2890):

then give the choices

OpenStudy (chriss):

3 + (1 + 11) = 15

OpenStudy (neer2890):

@ChrisS is right because associative property is (a+b)+c=a+(b+c)

OpenStudy (anonymous):

ty

OpenStudy (neer2890):

:)

OpenStudy (anonymous):

Here is an addition sentence. (14 + 13) + 2 = 29 Choose an equivalent addition sentence that shows the associative property of addition. 29 = (14 + 13) + 2 (14 + 13) + 2 = 29 14 + (13 + 2) = 29 14 + 13 + 2 = 29

OpenStudy (chriss):

yw

OpenStudy (neer2890):

use same method as above

OpenStudy (anonymous):

buhh i still dnt get it

OpenStudy (neer2890):

ok (14+13)+2=14+(13+2)=29 got it?

OpenStudy (anonymous):

yes now i get it

OpenStudy (neer2890):

good...:)

OpenStudy (neer2890):

so what's the answer?

OpenStudy (anonymous):

what u mean wats the new answer?

OpenStudy (anonymous):

o nvm its C

OpenStudy (neer2890):

yeah...C is correct

OpenStudy (anonymous):

kk ty

OpenStudy (neer2890):

you're welcome..:)

OpenStudy (anonymous):

The Tigers defeated the Dodgers in a baseball game that lasted 9 innings. By how many runs were the Dodgers defeated? Use the scorecard to help you answer the question. 2 runs 3 runs 4 runs 5 runs

OpenStudy (neer2890):

sorry...idk

OpenStudy (anonymous):

kk